# VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning

### v2

@@ I Introduction
However, there are still no VLN systems that can fully address the three capabilities. Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [12]. Classical approaches [9, 10, 29, 33], while efficient, struggle to integrate vision-language features. End-to-end learning methods [17, 24, 7, 26, 22] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios. Modular learning approaches [3, 21, 13, 5, 4], on the other hand, demonstrate robust real-world performance [12]. However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning. The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches. For instance, Vision-Language Frontier Maps (VLFM) [36] utilized VLMs to extract language-grounded features directly from RGB images. This enables the creation of a semantic map, guiding exploration based on human semantic knowledge. However, their reliance on computationally intensive models limits their deployability on low-power platforms. Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained vision-language cues.
@@ II Related Works
Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches. The Vision-Language Frontier Maps (VLFM) method [36] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration. VLFM achieves SOTA performance in the simulated environment, addressing the limitations of prior frontier-based and goal-oriented modular methods [33, 11, 3, 15]. However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN. Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for most real-world applications.

### v3

@@ I Introduction
However, there are still no VLN systems that can fully address the three capabilities. Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [13]. Classical approaches [9, 10, 31, 35], while efficient, struggle to integrate vision-language (VL) features. End-to-end learning methods [19, 26, 7, 28, 24] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios. Modular learning approaches [3, 23, 14, 5, 4], on the other hand, demonstrate robust real-world performance [13]. However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning. The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches [12, 38, 16, 39]. For instance, Vision-Language Frontier Maps (VLFM) [39] utilized VLMs to extract language-grounded features directly from RGB images. This enables the creation of a semantic map, guiding exploration based on human semantic knowledge. However, their reliance on computationally intensive models limits their deployability on low-power platforms. Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained VL cues.
@@ II Related Works
Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches [12, 38, 16, 39]. The Vision-Language Frontier Maps (VLFM) method [39] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration. VLFM achieves SOTA performance in the simulated environment, addressing the limitations of prior frontier-based and goal-oriented modular methods [35, 11, 3, 17]. However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN. Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for mobile-robot applications.